OpenStudy (anonymous):

use synthetic division and the remainder theorem to find p(a). p(x)=x^2+7x^2+4x. a= -2

OpenStudy (epikrebel):

\[p(x)=x^2+7x^2+4x\]?

OpenStudy (anonymous):

an equation can't be the answer. it has to be a number.

OpenStudy (epikrebel):

im asking its x^2+7x^2?

OpenStudy (anonymous):

woops, i accidently made x squared. it's P(x)=x^3+7x^2+4x; a= -2

OpenStudy (epikrebel):

When you say a=-2 you mean f(-2)?

OpenStudy (anonymous):

i have no idea what that means either. i guess it's f(-2). in the question it says, p(a) so it must be p(x)... i don't know, that's why i'm asking haha.

OpenStudy (anonymous):

yeah solve f(-2)

OpenStudy (epikrebel):

have no clue how to show synthetic division...

OpenStudy (anonymous):

well, just the answer would be fine. i just need to get this question correct. i've done fine on the other ones. this is just confusing.

OpenStudy (epikrebel):

Basically from synthetic division you should get 1 5 -6 12. Which is \[x^2+5x-6\] with the remainder of 12. And f(-2)=12 basically.

OpenStudy (epikrebel):

\[x^2+5x-6+\frac{12}{x+2}\]

OpenStudy (anonymous):

thank you.